In the coordinate plane the point X(0,1) is translated to the point X(5,4) under the same translation the points Y(-4,-3) and Z(3,5) are translated to Y and Z. What are the coordinates of Y and Z
1
Expert's answer
2020-02-27T12:15:57-0500

The translation vector is:

"T = X2 - X1 = (5-0, 4-1) = (5, 3).\\"

To obtain now points one should add this vector to the initial points:

"Y2 = Y1 + T = (-4 + 5, -3 + 3) = (1, 0)\\
Z2 = Z1 + T = (3 + 5, 5 + 3) = (8, 8)\\"

Answer:

"(1,0), (8,8)"